One radio transmitters operating at 64.5 MHz is 9.
Answer #1

wavelength = λ = c/f = 3x10⁸ / 64.5x10⁶ = 4.65m

AB is exactly 2λ so a standing wave, length 2λ, is set-up between A and B,

Consider what happens at each end, The waves will meet out of phase, so there will be a node at each end. Node to antinode distance = λ/4. The pattern of nodes (A) and antinodes (N) will be
N-A-N-A-N-A-N-A-N

Antinodes are where the 2 beams are in phase.

So from end A we must move a distance λ/4 to get to the nearest point where the two beams are in phase (the first antinode). So the answer is 4.65/4 = 1.1625m.